Archive for the Apiary Meetings Category

Apiary Meeting 15 October 2017

Weather unusually warm >20 degrees.  Bees flying and bringing in pollen.

Acetic acid treatment completed and pads removed.  Entrance blocks out in place but due to warm weather and high activity levels, mouseguards not yet placed.  Woodpecker guards put around hives.   Under-supered hives 2, 2A, 4, 7 & 8, i.e. all hives except 14×12 or double brood.

Last apiary meeting of the year.

Apiary Meeting Friday 6 October

Apistan strips removed.  Acetic acid treatment on brood boxes and supers.  Checked for stores and replenished where necessary.

Apiary Meeting 20 August 2017

Weather 20 degrees and cloudy.  Varroa treatment (Apistan) undertaken on all hives, including new beekeepers’ hives.
Batch 1609203035.  Expiry 10/2019.

Hive #1 – inspected by Richard.  Queen seen, still with damaged abdomen but continuing to lay.  Brood on 5 frames.  Some deformed wing virus and wax moth.  One wet super removed. Varroa = 50.

Hive #2 inspected by ?  Three emerged queen cells. No young brood.  Knocked down 4 queen cells.  Wet super not cleared and being re-filled.  One super of stores.

Hive # 2A – inspected by Mel.  2 x brood boxes.  Apistan applied to bottom brood only – brood on 6 frames.  Queen seen in bottom brood. Some sac brood.

Hive #3 – inspected by Mel.  2x deeps.  Queen not seen.  Brood on 4 frames.  Plenty of stores in top deep.  Varroa = 32.

Hive #4 – inspected by Anne.  Queen seen.  Brood on 6 frames. One capped queen cell. One uncapped queen cell in supercedure position (did not remove).  One uncapped queen cell removed.  4 frames not drawn, one moved to encourage drawing.

Hive # 6 – inspected by Anne.  2x deeps. Queen not seen. Brood in top box.

Hive #7 – inspected by Brad.  Queen not seen. Evidence of bald brood, chalk brood and parasitic mite syndrome.

Hive #8 – inspected by Liz.  Queen not seen.  Wet super being re-filled. Moved frames with stores below crown board.  Varroa = 29.

Observation hive – inspected by Rowan. Queen seen. Brood in 3 frames. Evidence of sac brood.  Suggest may not be suitable,Emma’s observation hive.

 

 

 

 

 

 

 

 

 

 

 

 

 

Apiary Record Sunday 14th May

Breezy, 18C  David Parkinson making Nuc Hives and initiating queen rearing.

Hive #1  Inspected by Richard Williams.  Good temper.   14×12 +shallow.  2 supers.   4 of 5 QC’s removed.  Sealed brood only. Sufficient space and stores.

Hive #3  Inspected by Liz Hawes.  Good temper.  2 deep.  3 supers.  Queen seen, not marked. 2 QC’s removed.  Sufficient space .

Hive #5  Inspected by Anne-Chantal Ballard.  Good temper. 14×12.  2 supers.  9 brood, eggs, 4 QC’s. Shallow added to brood.  QC’s & 1 framebrood removed to nucs.

 

Apiary Record 10th May

Sunny 20C

Hive #2  Inspected by Tom Blackburn  Artificial Swarm performed to create Hive #2a.

Hive #2 1 deep 1 super , queen marked blue, 2brood, 2 stores.

Hive #2a  2 deep, 2 sealed QC’s.Sufficient stores in deep

Hive #6  Inspected by Tom Blackburn.  2 deep, 6 & 7.5 frames brood 3 brood frames stores;  2 supers with limited stores.

Apiary Record Sunday 7th May

Sunny 17C .

Hive #1 Inspected by Liz Hawes.  14×12 + shallow.  2 supers.  Sufficient room & stores . Good Temper.  Queen not seen, swarm suspected. QC’s: 1 hatched, 2 torn down, 3 small, 2 good remaining.

Hive # 3  Inspected by Rowan Roberts.  2 deeps. 2 supers.  Sufficient room & stores . moderate Temper.  Queen not seen. 1 Queen cell removed to Nuc. hive, others destroyed.

Hive #4  Not inspected

Hive #5   Inspected by Rowan Roberts.  14×12. 2 supers.  Sufficient room & stores . moderate Temper.  Queen seen. 9 frames brood.  Bias.  Drone trap moved to brood edge.

Hive #7  Not inspected

Hive #8  Inspected by Phil Gurney. 1deep + shallow. 1 super.  good temper. 9 deep, 1 shallow brood, 1 torn down QC.  no eggs seen, young larvae.  Stores in brood only.

Hive # 9    Inspected by Phil Gurney. 1 deep . 1 super.  good temper.  No eggs, little larvae.  Poor condition, Sac brood.  Little stores, little feed taken.  Graham Rowden took Decision to destroy

 

 

Apiary Record 3rd May

Cloudy, cold & turning wet.  Quick inspection of Hive#2; very brief of hive#6 Tom Blackburn leading.

Hive #2    2 deeps with 4 brood in each.  Unsealed QC removed. 1 super with 9.5 stores.  Space.

Hive #6  Deteriorating weather, bad tempered bees.  Stores only inspected, 4 frames.

 

Wednesday 17th May Apiary meeting

Wednesday 17th May Apiary meeting cancelled due to bad weather

Apiary Meeting 30 April

Too cold for general inspections – 10 degrees. Demonstration of Queen marking using workers found in supers. Checks on nucs created the previous week.

Hive #4 – inspected by David P. 3 supercedure cells, likely to be emergency supercedures. Two cells cut from hive and pasted into Nuc 4A. Bees from some frames shaken into Nuc. Leave for 4 weeks. Next inspection 28 May.

Hive #3A – inspected by David P. Nuc created 23 April. Queenless.

Hive 4A – inspected by David P. Two supercedure cells pasted into Nuc from #4. Bees shaken into Nuc from some #4 frames.

Apiary Meeting 26 April

Brief inspections on Hives 2 & 6 only. Weather cold and wet.

Hive #2 – inspected by Liz & Rowan. Deep x 2. Brood on 4 frames. Stores on 5 frames. Some play cups. Stores in tip deep but not in super.

Hive #6 – inspected by Liz & Rowan. Deep x 2. Play cups seen in top deep. Only 5 frames inspected in bottom deep – hail storm prevented further checks.